# Parafermions with symmetry-protected non-Abelian statistics

### 14 Mar 2024 v1

@@ II.3 Effective braiding Hamiltonian
[…]
@@
where $a_{m}^{*}=a_{N-m}$, $b_{n}^{*}=b_{N-n}$, $c_{k,l}^{*}=c_{N-k,N-l}$ for hermitianity. The first and second terms in Eq. (8) denote the exchange Hamiltonian of two edges. The third term in $H_{E}$ couples $\alpha_{i}$-sector and $\beta_{i}$-sector and indicates that, generally speaking, time-periodic dynamics involving two pairs of PZMs cannot be divided into two independent processes. Thus to prove the SPNA statistics, we then explore certain conditions that are specific to braiding process to eliminate the third term in Eq. (8).
The braiding of two pairs of PZMs should satisfy anyon theory and spin-statistics theorem. The parity conservation of each sector suggests the following fusion rules of unitary symmetry-protected PZMs
@@
where $a\oplus b=(a+b)\text{ mod }N$, $F$ denotes a domain wall with a pair of PZMs, or equivalently, a single zero-energy Fock parafermion mode FockPara and $\psi_{q_{\alpha},q_{\beta}}$ denotes the Abelian anyon with $q_{\alpha(\beta)}$ quasiparticles of $\alpha_{i}(\beta_{i})$-sector. Since the Abelian anyon $\psi_{q_{\alpha},q_{\beta}}$ is a composite object, its topological spin $\theta_{q_{\alpha},q_{\beta}}$ can be deduced from those of $\psi_{1,0}$, $\psi_{0,1}$ and $\psi_{1,1}$. Suppose that the topological spins of $\psi_{1,0}$, $\psi_{0,1}$, $\psi_{1,1}$ are $\theta_{1,0}=\exp(i\phi_{\alpha})$, $\theta_{0,1}=\exp(i\phi_{\beta})$, $\theta_{1,1}=\exp[i(\phi_{\alpha}+\phi_{\beta}+\varphi)]$ separatedly, then the general expression for topological spin is $\theta_{q_{\alpha},q_{\beta}}=\exp[i(q_{\alpha}^{2}\phi_{\alpha}+q_{\beta}^{2}\phi_{\beta}+q_{\alpha}q_{\beta}\varphi)]$ as shown in Fig. 2 where $\theta_{2,1}$ is depicted as an example. The $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ structure of $\theta_{q_{\alpha},q_{\beta}}$ enforced by the fusion rules further requires $\theta_{q_{\alpha}+N,q_{\beta}}=\theta_{q_{\alpha},q_{\beta}}$ and $\theta_{q_{\alpha},q_{\beta}+N}=\theta_{q_{\alpha},q_{\beta}}$, which give
@@
where $u$, $v$, $w$ are integer numbers. Label the eigenstate of the system $|q_{\alpha},q_{\beta}\rangle$ by the presence of Abelian anyon $\psi_{q_{\alpha},q_{\beta}}$, then the braiding operator $U(T)$ takes a matrix form with only diagonal matrix element $\left[U(T)\right]_{q_{\alpha},q_{\beta}}=\langle q_{\alpha},q_{\beta}|U(T)|q_{\alpha},q_{\beta}\rangle$. From the spin-statistics theorem $\theta_{q_{\alpha},q_{\beta}}=\theta_{F}^{2}\left[U(T)\right]_{q_{\alpha},q_{\beta}}^{2}$ Kitaev2006 , where $\theta_{F}$ denotes the topological spin of a Fock parafermion mode, we find
$\left[U(T)\right]_{q_{\alpha},q_{\beta}}=(-1)^{f(q_{\alpha},q_{\beta})}e^{i\pi(q_{\alpha}^{2}u+q_{\beta}^{2}v+2q_{\alpha}q_{\beta}w)/2N},$ (12)
where ${f(q_{\alpha},q_{\beta})}$ is an integer-valued function. This formula indicates that the matrix elements of $U(T)$ can only take discrete values and $\left[U(T)\right]^{4N}=1$. From Eq (12), we can also read the diagonal matrix element of the effective braiding Hamiltonian $\langle q_{\alpha},q_{\beta}|H_{E}|q_{\alpha},q_{\beta}\rangle=(q_{\alpha}^{2}u+q_{\beta}^{2}v+2q_{\alpha}q_{\beta}w)[\pi/(2NT)]+(\pi/T)f(q_{\alpha},q_{\beta})$, thus the originally continuous coefficients $a_{m}$, $b_{n}$, $c_{k,l}$ in Eq. (8) should be similarly restricted to discrete values.
The symmetric effective braiding Hamiltonian Eq. (8) and the property $\left[U(T)\right]^{4N}=1$ are sufficient to the proof to unitary SPNA statistics of PZMs. The property $\left[U(T)\right]^{4N}=1$ or equivalently the discrete values of the matrix elements of $U(T)$ imply that the coefficients $(a_{m},b_{n},c_{k,l})$ in the effective braiding Hamiltonian must be discrete values which cannot change continuously. Thus the specific values of the coefficients can be determined by a special case as long as the Hamiltonian in the special case is topologically equivalent to that in the generic case. Futhermore, in a special case where the two sectors of PZMs $\alpha_{1,2}$ and $\beta_{1,2}$ belong to two fully decoupled subsystem copies, we must have $c_{k,l}=0$ for all the $k$, $l$. Thus the braiding operator should also be decoupled for the generic case where the whole system cannot be decoupled into two independent subsystems. In this case, the PZMs are braided independently in each sector. Then the braiding operator reads
@@
where $\lambda_{\alpha(\beta),m}=\omega^{m(m+N+2r_{\alpha(\beta)})/2}$ with $r_{\alpha(\beta)}=0,\cdots,N-1$. This formula is the main result of this work and gives the first example of SPNA statistics in correlated systems. Our proof only relys on the unitary symmetry protection and the general property of braiding process. When $N=2$, the Eq. (13) is reduced to the braiding operator of two symmetry-protected Majorana doublets XJL2014PRX ; Gao2016 ; JSH2022 or two Dirac zero modes Loss2013fractional ; YWu2020a ; YWu2020b ; YWu2022 .
@@ III Braiding matrix in a tri-junction
We can also establish the SPNA statistics by explicitly deriving the non-Abelian geometric phases resulting from the manipulation of PZMs within a tri-junction configuration. The exchange of PZMs is accomplished through a three-point turn in the tri-junction, with the groundstate degeneracy remaining invariant throughout the adiabatic exchange process. In the following subsections, we expound upon the braiding dynamics within the tri-junction and delve into the characteristics of degenerate groundstates of the system. Subsequently, we derive the recursion relation for non-Abelian Berry phases accumulated during the braiding process. Finally, we demonstrate that the Berry phases necessarily adopt a decoupled form, as dictated by the adiabatic condition.
@@ III.1 Tri-junction and groundstate subspace
The braiding process consists of three subsequent stages of adiabatic time evolution. Consider six pairs of symmetry-protected PZMs $\alpha_{1,\cdots,6}$ and $\beta_{1,\cdots,6}$ as shown in Fig. 3(a) and suppose that we are going to braid $\alpha_{1}$, $\beta_{1}$ and $\alpha_{2}$, $\beta_{2}$. Before the braiding, we initialize the system by coupling $\alpha_{5}$, $\beta_{5}$ and $\alpha_{6}$, $\beta_{6}$ together, then only $\alpha_{1,\cdots,4}$ and $\beta_{1,\cdots,4}$ remain zero-energy. Since the total parity for each sector is conserved for symmetry-protected parafermionic systems during the braiding, if the parity for $\alpha(\beta)$-sector is fixed, then the groundstates after initialization of this system is $N\times N$-fold degenerate. At the first stage of braiding process, the coupling between $\alpha_{1}$, $\beta_{1}$ and $\alpha_{6}$, $\beta_{6}$ is turned on adiabatically, and the coupling between $\alpha_{5}$, $\beta_{5}$ and $\alpha_{6}$, $\beta_{6}$ is turned off adiabatically at the same time. The coupling Hamiltonian for this stage is written as $H_{\text{I}}(t)=\eta_{16}(t)H_{16}+\eta_{56}(t)H_{56}$, where $\eta_{16}(t)$, $\eta_{56}(t)$ are depicted in Fig. 3(c1) and $H_{ij}$ denotes the coupling term between $\alpha_{i}$, $\beta_{i}$ and $\alpha_{j}$, $\beta_{j}$, whose concrete form will be specified later. The second and third stages of the braiding process are governing by similar Hamiltonians $H_{\text{II}}(t)=\eta_{26}(t)H_{26}+\eta_{16}(t)H_{16}$ and $H_{\text{III}}(t)=\eta_{56}(t)H_{56}+\eta_{26}(t)H_{26}$ as depicted in Fig 3. After all the three stages, the Hamiltonian of the whole system is returned to the original form after initialization and this braiding process is finished. During the braiding, there are $N\times N$ degenerate groundstates, and since $\alpha_{3}$, $\beta_{3}$ and $\alpha_{4}$, $\beta_{4}$ are zero-modes at each stage, the instantaneous goundstates can always be labelled as the parity eigenstates of $\alpha_{3}$, $\alpha_{4}$ and $\beta_{3}$, $\beta_{4}$
@@
where $Q_{X,34}=\omega^{(N+1)/2}X_{3}X_{4}^{\dagger}$ with $X=\alpha,\beta$, $q^{\prime}_{X}$ is $q_{X,34}$ for short. In these bases, the effect of braiding is manifested in the Berry phase of the groundstate $|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{f})\rangle=e^{i\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{i})\rangle$, where $t_{i(f)}$ denotes the initial (final) time of the braiding dynamics.
@@ III.2 Recursion relation for Berry phases
From the above recursion relations, we see that if the phase factor $\delta_{X,i(f)}(q^{\prime}_{\alpha},q^{\prime}_{\beta})$ can be divided into two independent parts $\delta_{X,i(f)}(q^{\prime}_{\alpha},q^{\prime}_{\beta})=\delta^{\alpha}_{X,i(f)}(q^{\prime}_{\alpha})+\delta^{\beta}_{X,i(f)}(q^{\prime}_{\beta})$, where $\delta^{\alpha}_{X,i(f)}(\delta^{\beta}_{X,i(f)})$ only relies on a single parity $q^{\prime}_{\alpha}(q^{\prime}_{\beta})$, then the total Berry phase $\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})$ is also decoupled.
@@ III.3 Decoupled braiding matrix
The adiabatic condition implies the decoupled forms of the phase factor $\delta_{X,i(f)}$ and the non-Abelian Berry phases. To see this, we explore the concrete form of the coupling Hamiltonian $H_{ij}$. Most generically,
@@
Certain Hamiltonians like $H_{ij}^{\alpha}=H_{ij}^{\beta}=0$, $H_{ij}^{\text{int}}=c^{\prime}_{1,1}Q_{\alpha,ij}Q_{\beta,ij}+h.c.$ are forbidden since their groundstates are degenerate. Then the physical coupling Hamiltonian must be continuously connected to the simple one $H^{\prime}_{ij}=a^{\prime}_{1}Q_{\alpha,ij}+b^{\prime}_{1}Q_{\beta,ij}+h.c.$, whose groundstate is non-degenerate. The Hamiltonian $H^{\prime}_{ij}$ takes a decoupled form, as a result, the braiding matrix can be divided into two independent matrices for the two sectors of PZMs
$\gamma(q^{\prime}_{\alpha},q^{\prime}_{\beta})=\frac{\pi}{N}\left[(q^{\prime}_{\alpha}-k_{\alpha})^{2}+(q^{\prime}_{\beta}-k_{\beta})^{2}\right],$ (22)
@@ IV Symmetry-protected PZMs in interacting quantum wires
We provide a nanowire model hosting a pair of PZMs at each edge protected by mirror symmetry. In this model, the PZMs are naturally classified into two sectors with conserved parity. However, symmetry-preserving interactions can mix the PZMs of the two sectors without ruining non-Abelian statistics, which serves as a non-trivial demonstration of our generic theory.

### 21 Mar 2024 v2

@@ II.3 Effective braiding Hamiltonian
We now turn to studying the braiding of PZMs through an effective braiding Hamiltonian under the unitary symmetry protection. By definition, the braiding process is an adiabatic evolution $U(T)=\mathcal{T}\exp\bigr{[}{-i\int_{-T/2}^{T/2}H(\tau)d\tau}\bigr{]}$, with the Hamiltonian returning to initial one after braiding $H(-T/2)=H(T/2)$ and $\mathcal{T}$ the time-ordering operator, and can be described by an effective Floquet Hamiltonian $H_{E}=(i/T)\ln{U(T)}$. If the unitary symmetry preserves instantaneously, such that $[H(\tau),S]=0$ for every moment within the interval $-T/2<\tau<T/2$, then the effective Hamiltonian also satisfies the symmetry $[H_{E},S]=0$ as the unitary time evolution operator $U(T)$ commutes with unitary symmetry. Consequently, the effective Hamiltonian preserves the parity of each sector, and it must take the generic form as
@@
where $a_{m}^{*}=a_{N-m}$, $b_{n}^{*}=b_{N-n}$, $c_{k,l}^{*}=c_{N-k,N-l}$ for hermitianity. The first and second terms in Eq. (10) denote the exchange Hamiltonian of two sectors. The third term in $H_{E}$ couples $\alpha_{i}$-sector and $\beta_{i}$-sector and indicates that, with only parity conservation, time-periodic dynamics involving two pairs of PZMs cannot be divided into two independent processes. Thus to prove the SPNA statistics, we shall show that the third term in Eq. (10) should indeed be absent in the braiding.
The braiding of two pairs of PZMs should satisfy anyon theory and spin-statistics theorem. The parity conservation of each sector leads to the following fusion rules of the unitary symmetry-protected PZMs
@@
where $a\oplus b=(a+b)\text{ mod }N$, $F$ denotes a domain wall with a pair of PZMs, or equivalently, a single zero-energy Fock parafermion mode FockPara and $\psi_{q_{\alpha},q_{\beta}}$ denotes the Abelian anyon with $q_{\alpha(\beta)}$ quasiparticles of $\alpha_{i}(\beta_{i})$-sector. Since the Abelian anyon $\psi_{q_{\alpha},q_{\beta}}$ is a composite object, its topological spin $\theta_{q_{\alpha},q_{\beta}}$ can be deduced from those of $\psi_{1,0}$, $\psi_{0,1}$ and $\psi_{1,1}$. Suppose that the topological spins of $\psi_{1,0}$, $\psi_{0,1}$, $\psi_{1,1}$ are $\theta_{1,0}=\exp(i\phi_{\alpha})$, $\theta_{0,1}=\exp(i\phi_{\beta})$, $\theta_{1,1}=\exp[i(\phi_{\alpha}+\phi_{\beta}+\varphi)]$ separately. Then the general expression for topological spin is $\theta_{q_{\alpha},q_{\beta}}=\exp[i(q_{\alpha}^{2}\phi_{\alpha}+q_{\beta}^{2}\phi_{\beta}+q_{\alpha}q_{\beta}\varphi)]$ as shown in Fig. 2 where $\theta_{2,1}$ is depicted as an example. The $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ structure of $\theta_{q_{\alpha},q_{\beta}}$ enforced by the fusion rules further requires $\theta_{q_{\alpha}+N,q_{\beta}}=\theta_{q_{\alpha},q_{\beta}}$ and $\theta_{q_{\alpha},q_{\beta}+N}=\theta_{q_{\alpha},q_{\beta}}$, which give
@@
where $u$, $v$, $w$ are integer numbers. Label the eigenstate of the system $|q_{\alpha},q_{\beta}\rangle$ by the presence of Abelian anyon $\psi_{q_{\alpha},q_{\beta}}$, in which the braiding $U(T)$ takes a matrix form with only diagonal matrix elements $\left[U(T)\right]_{q_{\alpha},q_{\beta}}=\langle q_{\alpha},q_{\beta}|U(T)|q_{\alpha},q_{\beta}\rangle$. From the spin-statistics theorem $\theta_{q_{\alpha},q_{\beta}}=\theta_{F}^{2}\left[U(T)\right]_{q_{\alpha},q_{\beta}}^{2}$ Kitaev2006 , where $\theta_{F}$ denotes the topological spin of a Fock parafermion mode, we find
$\left[U(T)\right]_{q_{\alpha},q_{\beta}}=(-1)^{f(q_{\alpha},q_{\beta})}e^{i\pi(q_{\alpha}^{2}u+q_{\beta}^{2}v+2q_{\alpha}q_{\beta}w)/2N}.$ (14)
Here ${f(q_{\alpha},q_{\beta})}$ is an integer-valued function. This formula tells that the elements of $U(T)$ can only take discrete values and $\left[U(T)\right]^{4N}=1$. From the Eq (14), we can further obtain the diagonal matrix element of the effective braiding Hamiltonian $\langle q_{\alpha},q_{\beta}|H_{E}|q_{\alpha},q_{\beta}\rangle=(q_{\alpha}^{2}u+q_{\beta}^{2}v+2q_{\alpha}q_{\beta}w)[\pi/(2NT)]+(\pi/T)f(q_{\alpha},q_{\beta})$, which determines that the coefficients $a_{m}$, $b_{n}$, $c_{k,l}$ in Eq. (10) must also take discrete values.
The symmetry of the braiding Hamiltonian in Eq. (10) and the discreteness of $U(T)$ suffice to prove the unitary SPNA statistics of PZMs. The discreteness of the matrix elements in $U(T)$ implies that the coefficients $(a_{m},b_{n},c_{k,l})$ in the braiding Hamiltonian cannot change continuously. This key feature facilitates to determine the coefficients from a special case that the whole system hosting PZMs is formed by two independent subsystem. The generic system can be obtained from such special case by adiabatically switching on couplings and interactions between the subsystems without undergoing topological phase transition. Then the braiding matrix obtained in the special case must be applicable to the generic case. In a special case that the PZMs $\alpha_{1,2}$ and $\beta_{1,2}$ belong to two fully decoupled subsystems, we must have that $c_{k,l}=0$ for all $k$ and $l$. With this we conclude that the generic braiding operator reads
@@
where $\lambda_{\alpha(\beta),m}=\omega^{m(m+N+2r_{\alpha(\beta)})/2}$ with $r_{\alpha(\beta)}=0,\cdots,N-1$. The above result shows that the braiding of the PZM pairs with unitary symmetry protection can always be reduced to braiding the two copies of PZMs individually, even though the whole system cannot be characterized by two decoupled subsystems due to the generally existing strong interactions. This braiding operation is highly nonlinear for $N>2$, sharply contrast to the Majorana counterparts with symmetry protection XJL2014PRX ; Gao2016 ; JSH2022 , in which the braiding transformation is linear.
@@ III Braiding matrix in a tri-junction
To further confirm the SPNA statistics of PZMs, we study the braiding dynamics by explicitly deriving the non-Abelian geometric phases resulting from adiabatic manipulation of PZMs within a tri-junction configuration in Fig. 3. The exchange of PZMs is accomplished through a three-point turn in the tri-junction, with the groundstate degeneracy remaining invariant throughout the adiabatic exchange process. In the following subsections, we expound upon the braiding dynamics within the tri-junction and delve into the characteristics of degenerate groundstates of the system. Subsequently, we derive the recursion relation for non-Abelian Berry phases accumulated during the braiding process. Finally, we show that the Berry phases necessarily adopt a symmetric form, as dictated by the adiabatic condition.
@@ III.1 Tri-junction and groundstate subspace
The braiding process consists of three subsequent stages of the adiabatic time evolution. Consider six pairs of symmetry-protected PZMs $\alpha_{1,\cdots,6}$ and $\beta_{1,\cdots,6}$ as shown in Fig. 3(a) and suppose that we are going to braid $\alpha_{1}$, $\beta_{1}$ and $\alpha_{2}$, $\beta_{2}$. Before the braiding, we initialize the system by coupling $\alpha_{5}$, $\beta_{5}$ and $\alpha_{6}$, $\beta_{6}$ together, then only $\alpha_{1,\cdots,4}$ and $\beta_{1,\cdots,4}$ remain zero-energy. Since the total parity for each sector is conserved for symmetry-protected parafermionic systems during the braiding, if the parity for $\alpha(\beta)$-sector is fixed, then the groundstates after initialization of this system is $N\times N$-fold degenerate. At the first stage of braiding process, the coupling between $\alpha_{1}$, $\beta_{1}$ and $\alpha_{6}$, $\beta_{6}$ is turned on adiabatically, and the coupling between $\alpha_{5}$, $\beta_{5}$ and $\alpha_{6}$, $\beta_{6}$ is turned off adiabatically at the same time. The coupling Hamiltonian for this stage is written as $H_{\text{I}}(t)=\eta_{16}(t)H_{16}+\eta_{56}(t)H_{56}$, where $\eta_{16}(t)$, $\eta_{56}(t)$ are depicted in Fig. 3(c1) and $H_{ij}$ denotes the coupling term between $\alpha_{i}$, $\beta_{i}$ and $\alpha_{j}$, $\beta_{j}$, whose concrete form will be specified later. The second and third stages of the braiding process are governing by similar Hamiltonians $H_{\text{II}}(t)=\eta_{26}(t)H_{26}+\eta_{16}(t)H_{16}$ and $H_{\text{III}}(t)=\eta_{56}(t)H_{56}+\eta_{26}(t)H_{26}$ as depicted in Fig 3. After all the three stages, the Hamiltonian of the whole system returns to the original form after initialization and this braiding process is completed. During the braiding, there are $N\times N$ degenerate groundstates, and since $\alpha_{3}$, $\beta_{3}$ and $\alpha_{4}$, $\beta_{4}$ are zero-modes at each stage, the instantaneous goundstates can always be labelled as the parity eigenstates of $\alpha_{3}$, $\alpha_{4}$ and $\beta_{3}$, $\beta_{4}$
@@
where $Q_{X,34}=\omega^{(N+1)/2}X_{3}X_{4}^{\dagger}$ with $X=\alpha,\beta$, and $q^{\prime}_{X}$ is $q_{X,34}$ for short. In these bases, the effect of braiding is manifested in the Berry phase of the groundstate $|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{f})\rangle=e^{i\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{i})\rangle$, where $t_{i(f)}$ denotes the initial (final) time of the braiding dynamics.
@@ III.2 Recursion relation for Berry phases
From the above recursion relations, we see that if the phase factor $\delta_{X,i(f)}(q^{\prime}_{\alpha},q^{\prime}_{\beta})$ can be divided into two independent parts $\delta_{X,i(f)}(q^{\prime}_{\alpha},q^{\prime}_{\beta})=\delta^{\alpha}_{X,i(f)}(q^{\prime}_{\alpha})+\delta^{\beta}_{X,i(f)}(q^{\prime}_{\beta})$, where $\delta^{\alpha}_{X,i(f)}(\delta^{\beta}_{X,i(f)})$ only relies on the parity $q^{\prime}_{\alpha}(q^{\prime}_{\beta})$, the total Berry phase $\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})$ can be similarly decomposed.
@@ III.3 Decoupled braiding matrix
The adiabatic condition determines the decomposition of the phase factor $\delta_{X,i(f)}$ and the non-Abelian Berry phases. To see this, we explore the concrete form of the coupling Hamiltonian $H_{ij}$. Most generically,
@@
Certain Hamiltonians like $H_{ij}^{\alpha}=H_{ij}^{\beta}=0$, $H_{ij}^{\text{int}}=c^{\prime}_{1,1}Q_{\alpha,ij}Q_{\beta,ij}+h.c.$ are forbidden since their groundstates are degenerate. Then the physical coupling Hamiltonian must be continuously connected to the simple one $H^{\prime}_{ij}=a^{\prime}_{1}Q_{\alpha,ij}+b^{\prime}_{1}Q_{\beta,ij}+h.c.$, whose groundstate is non-degenerate. The Hamiltonian $H^{\prime}_{ij}$ takes a decoupled form and, as a result, the braiding matrix can be divided into two independent matrices for the two sectors of PZMs with the matrix elements of the Berry phases
$\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})=\frac{\pi}{N}\left[(q^{\prime}_{\alpha}-k_{\alpha})^{2}+(q^{\prime}_{\beta}-k_{\beta})^{2}\right],$ (24)
@@ IV Symmetry-protected PZMs in interacting quantum wires
Finally we propose a physical system based on strongly correlated nanowires to realize a pair of PZMs at each edge protected by mirror symmetry. In this model, the PZMs are naturally classified into two symmetry sectors with conserved parity. Further, symmetry-preserving interactions can generally mix the PZMs of the two sectors without ruining the non-Abelian statistics, which serves as a non-trivial demonstration of the generic theory.